OpenStudy (anonymous):

The domain of the function f(x) = 3x3 is {2, 5}. What is the function’s range? The range of the function f(x) = 3x3 when its domain is {2, 5} is

OpenStudy (anonymous):

\[f(x)=3x^3\] right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i couldnt get it like that sorry

OpenStudy (anonymous):

if the domain is the two numbers \(2\) and \(5\) then the range is \[f(2)=3\times 2^3\] and \[f(5)=3\times 5^3\]

OpenStudy (anonymous):

of course they probably want you to actually compute those numbers

OpenStudy (anonymous):

btw both those questions are the same, just worded differently

OpenStudy (anonymous):

yea, im confused which one is the answer

OpenStudy (anonymous):

the answer is two numbers, not one compute them both

OpenStudy (anonymous):

they only have a answer blank for this one - The range of the function f(x) = 3x3 when its domain is {2, 5} is

OpenStudy (anonymous):

so it would be f(5)=3×5^3 ? right

OpenStudy (anonymous):

okay nvm

OpenStudy (anonymous):

two answers really

OpenStudy (anonymous):

maybe you are supposed to separate them by a comma or sommat